# Signatures of Enhanced Superconducting Properties in Niobium Cavities

### v3

@@ Abstract
Superconducting radio-frequency (SRF) niobium cavities are critical for modern particle accelerators, as well as for advancing superconducting quantum systems and enabling ultra-sensitive searches for new physics. In this work, we report a systematic observation of an anomalous frequency dip in Nb cavities, which occurs at temperatures just below the critical temperature ($T_{\mathrm{c}}$), indicative of enhanced superconducting properties at $T\ll T_{c}$. The magnitude of this dip is strongly correlated with the RF surface resistance, impurity distribution near the surface, and $T_{\mathrm{c}}$. We also present the first demonstration of the coherence peak in the AC conductivity of Nb SRF cavities. By comparing recent theories developed to model this experimental data, we show that the frequency dip feature, larger coherence peak height, and reduction in the temperature-dependent surface resistance with RF current occur at minimal but finite levels of disorder.
@@ I Introduction
Superconducting radio frequency (SRF) niobium cavities are record-high quality factor $Q_{0}>10^{10}-10^{11}$ man-made resonators that serve as the primary accelerating structures in modern particle accelerators [1], longest coherence microwave quantum superconducting systems [2], and an ultrasensitive platform for dark sector and dark matter candidate searches [3, 4, 5, 6]. These ultra-high quality factors are enabled by the low microwave surface resistance $R_{s}$ ($\propto 1/Q_{0}$) of superconducting Nb. Typically, $R_{s}$ increases with the magnitude of the peak surface magnetic field $B_{p}$ due to increased screening currents. Extensive research over the past decades has allowed the discovery and mitigation of several such field-dependent physical mechanisms in various regimes [1]. This most recently includes proximity-coupled niobium nanohydrides [7] and their mitigation by oxygen inward diffusion [8] and the reduction of two-level system-driven dissipation by oxide removal [9, 2].
@@
The AC complex conductivity enables further insight on quasiparticles and superconducting carriers. For frequencies far below the superconducting gap ($\hbar\omega\ll\Delta$), $\sigma_{1}$ exhibits the coherence peak, a nonmonotonic dependence of the quasiparticle conductivity with temperature that arises due to coherence factors in BCS theory [15] and is analogous to the Hebel-Slichter peak in nuclear spin relaxation [16]. However, until now, this coherence peak has not been reported in Nb SRF cavities. We show that the height of the coherence peak also correlates with the frequency dip in Fig.1.
In this article, a systematic study is presented on a large number ($>$40) of bulk niobium SRF cavities revealing the anomalous decrease (dip) in cavity resonant frequency at temperatures just below $T_{\mathrm{c}}$ in nitrogen-doped cavities. In addition, we report that the other four (non-dip) characteristic temperature dependencies uniquely correspond to other explored surface treatments, and thus correlate strongly with cavity performance in $R_{T}$ with the stored field. The measured coherence peak in the complex conductivity of bulk niobium SRF cavities is fitted with a phenomenological model incorporating a pair-breaking term in the BCS [15] density of states. We comment on different models proposed by other groups which help provide insight on the origins of these various phenomena. Fits with theories suggest that the coherence peak, along with the dip phenomenon, high $Q_{0}$, and $R_{T}$ decrease with field, are signatures of improved superconducting properties and occur when extrinsic pair-breaking scattering and pair-conserving scattering from nonmagnetic impurities are minimal but finite.
@@ II Experimental Procedure
A large set of bulk niobium SRF cavities of TESLA [17] elliptical shape were employed for our studies. The resonant frequencies of these cavities varied from 650 MHz to 3.9 GHz, which allowed probing the low frequency limit of Nb ($\hbar\omega\ll\Delta$). All cavities went through a bulk electropolishing of 120 $\mu$m material removal and 800C hydrogen degassing step [18] followed by various surface treatments. These treatments included $\sim$40 $\mu$m electropolishing (referred to as EP here), nitrogen infusion [19], 120C baking for 48 hours [20], 75/120C baking [21, 22] wherein cavities were first baked at 75C before the standard 120C bake, and nitrogen doping [10] with varying dopant concentrations. Dopant concentration is varied by altering the durations of baking in N atmosphere and in vacuum, which has the shorthand notation of minutes in N/minutes in vacuum (i.e., 2/6) and/or the amount of removal via EP.
All cavities were first cooled down to 2 K where standard measurements [23] of $Q_{0}$ as a function of the peak surface magnetic field $B_{\mathrm{peak}}$ have been performed. To extract $R_{T}$ from the measured $Q_{0}$, we used the methods laid out in [14].
@@
To calculate the AC conductivity, we used the method utilized by Trunin et al. [27]. To extract the surface resistance, we measured $Q_{0}$ from 1.5 K up to 10 K using a combination of standard power balance measurements in continuous wave and decay modes along with VNA measurements. We converted to surface resistance via $R_{s}=G/Q_{0}$, where $G=$ 270 $\Omega$. The surface reactance was measured utilizing the above discussed frequency vs temperature measurements from 6 K to 10 K and extrapolating to lower temperatures by means of fitting with the Halbritter routine [25] and using $X_{s}(T)=-2G\Delta f_{0}(T)/f_{0}+X_{n}$. We obtained the additive constant using the normal conducting surface resistance measured just above the transition temperature $R_{n}$(10 K). In the local limit, valid for dirty Nb, $R_{n}(\text{10 K})=X_{n}(\text{10 K})$. When the local limit is not applicable, as is true for high purity niobium [20], the correct relationship between $R_{s}$ and $X_{s}$ is obtained using the universal impedance curves plotted against the dimensionless parameter $\alpha$ calculated by Reuter and Sondheimer in the microwave region [28]. The parameter $\alpha$ is proportional to $\ell^{3}$/$\rho\ell$, where $\ell$ is the mean free path and $\rho\ell$ is a temperature independent material constant, which is 6 $\times$ 10-16 $\Omega$m2 for Nb [20]. Finally, the complex conductivity is calculated as
@@ III.3 Correlating $Q_{0}$ with $\Delta f_{dip}$
The evolution observed in Fig. 2 suggests correlation between the frequency dip and $Q_{0}$ ($\propto 1/R_{s}$); to investigate this, we plotted the quality factor measured at 68 mT against the dip magnitude for an ensemble of variously nitrogen doped cavities in Fig. 3. Indeed, we observe a linear trend, suggesting a potential relationship between the positive Q-slope and frequency dip phenomena. We note that the trend shown in Fig. 3 is relevant for the accelerator community as it facilitates a method for quickly estimating the quality factor of N-doped cavities at high fields from low field frequency domain measurements, which may be practical when high power infrastructure is not readily available.
Figure 3: Quality factor measured at 68 mT and a temperature of 2 K plotted against the dip magnitude for variously nitrogen doped cavities. Shown for reference are results obtained from a cavity post EP.
@@ III.5 Measurement of AC Complex Conductivity
To gain insight on the conditions under which different frequency features near $T_{c}$ occur, we measured the AC complex conductivity of two niobium cavities: i) the cavity presented in Fig. 2(a) with $\ell=550\pm 51$ nm which showed the standard feature near $T_{c}$ (which we call the EP cavity) and ii) and a cavity post nitrogen doping which exhibited the prominent dip. Fig. 5(a) and its inset shows the expected $Q_{0}(B_{p})$ and $R_{T}(B_{p})$ behaviors characteristic of cavities subjected to these surface treatment. Fig. 5(b) presents the real and the imaginary parts of the measure surface impedance of the two cavities. For the N-doped cavity, the transition temperature was 9.04 K; using Fig. 2(c), the extrapolated $\ell$ was about 70 nm.
@@
EP 3.70 $\times$ 10 -4 550 36 $\pm$ 4 40
N-Doped 5.88 $\times$ 10 -4 70 57 $\pm$ 6 48
@@
Figure 5: (a) Quality factor vs peak magnetic field of two 1.3 GHz cavities subjected to EP or N-doping. Inset shows the temperature dependent surface resistance at 2 K as a function of peak magnetic field. (b) Surface reactance of the EP and N-doped cavities as a function of temperature; inset depicts the surface resistance as a function of temperature. (c) Real ($\sigma_{1}$) and imaginary ($\sigma_{2}$) components of the AC conductivity normalized to the normal conducting value $\sigma_{n}$. Error bars for the test post N-doping are dominated by measurement uncertainty while those shown post EP are dominated by error in the fitted $\ell$ value. Dashed and solid lines show fits.
The calculated AC complex conductivities of the N-doped and EP cavities are presented in Fig. 5(c). We note that as the technique used to measure the surface impedance sums over the inner RF surface, these curves represent the average cavity response.
The non-monotonic dependence in the real part of the complex conductivity in BCS superconductors, called the coherence peak, arises due to the singularity in the superconducting density of states at the gap edge [15, 32]. According to Dynes et al. [33], inelastic scattering smears the singularity and introduces subgap quasiparticle states. By incorporating the Dynes smearing parameter $\Gamma$ into the MB conductivity, it is possible to obtain a phenolomenological measure of the pair-breaking processes present in a system by fitting the coherence peak using
@@
From Fig. 5(c), we observe that both the EP and N-doped cavities exhibit the so-called coherence peak, with the maxima at $\sim$0.9 $T/T_{c}$ differing in amplitude. This suggests a variation in the quasiparticle behavior within the RF layer due to distinct impurity distributions. Moreover, the stark difference observed in $\sigma_{2}/\sigma_{n}$ between the two cavities highlights a variance in the superconducting condensate behavior.
The dashed and solid curves in Fig. 5(c) are calculated using Eq. 2-4. We find that both $\sigma_{1}/\sigma_{n}$ and $\sigma_{2}/\sigma_{n}$ of the EP cavity are best modelled with an average superconducting gap $\Delta_{0}=1.5$ meV and an inelastic scattering parameter $\Gamma/\Delta_{0}$ = 0.025. The N-doped cavity is instead better fitted with a larger average superconducting gap $\Delta_{0}=1.6$ meV and a lower level of pair-breaking ($\Gamma/\Delta_{0}$ = 0.0066). Note that a larger coherence peak occurs for the cavity with greater elastic scattering (shorter $\ell$) due to the presence of N interstitial. According to Anderson’s theorem [34], elastic scattering off nonmagnetic impurities has no effect on superconductivity. As a result, while N impurities increase scattering, they likely mitigate depairing processes which results in a larger coherence peak.
@@ IV Discussion
The frequency phenomena and trends observed in this work have stimulated the development of new theories which seek to provide insight into their microscopic origins. Ueki et al. provided the first explanation of the anomalous near $T_{c}$ frequency phenomena [35] by calculating the complex surface impedance via Slater’s approach to solving Maxwell’s equations for the case of a cavity while considering anisotropy in the superconducting gap [30] and inhomogeneous disorder within the penetration depth. By varying the level of disorder within the cavity screening region, Ueki correctly calculated the foot, dip+bump, and dip features in the frequency response just before $T_{c}$. Moreover, the authors obtained the predicted cavity $Q_{0}$ as a function of disorder. We note that the frequency dip feature and a maximum in $Q_{0}$ were obtained for intermediate levels of disorder. Ueki also correctly reproduced the frequency dependence of the dip feature reported in Fig. 4. Zarea et al, instead, derived the eigenvalue equation for the fundamental TM mode of a cylindrical RF cavity which considers the current response obtained with the Keldysh formulation of the quasiclassical theory of superconductivity [36]. The result considers the confinement of the normal conducting and superconducting currents within the RF layer and shows that the anomalous frequency dip is a product of the competition between the normal metal skin depth and the London penetration depth. We note that Zarea also found that the frequency dip and a maximum in cavity $Q_{0}$ occur at intermediate levels of disorder.
@@
The results presented in Fig. 5 have also stimulated theoretical work by other groups which further improves our understanding of the mechanisms behind the improved performance of niobium SRF cavities in the presence of various near-surface impurity profiles. Herman and Hlubina [37] used the Nam approach [38] to calculate the optical response of a superconductor in the limit where energy tends to zero and with pair-conserving and pair-breaking elastic-scattering processes built into the energy and gap functions. The result describes the effect of subgap quasiparticle states on the optical conductivity and utilizes the concept of “Dynes superconductors.” The authors present excellent fits to the real and imaginary parts of the AC complex conductivity of both cavities presented in Fig. 5 and find that pair-breaking processes dominate the height of the coherence peak in $\sigma_{1}/\sigma_{n}$. Herman and Hlubina also highlight that the ratio between pair-breaking and pair-conserving scattering rates is key in determining the coherence peak extent and show that cleaner superconductors counterintuitively yield lower peak heights. Kubo, instead, used the Dynes $\Gamma$ in the Eilenberger formalism of the BCS theory to calculate several superconductor quantities in the presence of varying levels of nonmagnetic impurity scattering rates as a function of temperature and energy [39]. These calculations also well model the results presented in Fig. 5 and, similar to Herman and Hlubina, show that the ratio between pair-breaking and nonmagnetic impurity scattering rates is key in determining the height of the coherence peak. Moreover, Kubo finds that there exists a minimum in the surface resistance at finite levels of pair-breaking and pair-conserving scattering. We note that the phenomenologically derived Eqs. 2-4 in the present manuscript reproduce the salient feature of the above theories that greater smearing of the density of states yields a reduction of the coherence peak height and the zero-temperature value of $\sigma_{2}/\sigma_{n}$.

### v4

@@ Abstract
Superconducting radio-frequency (SRF) niobium cavities are critical for modern particle accelerators, as well as for advancing superconducting quantum systems and enabling ultra-sensitive searches for new physics. In this work, we report a systematic observation of an anomalous frequency dip in Nb cavities, which occurs at temperatures just below the critical temperature ($T_{\mathrm{c}}$), indicative of enhanced superconducting properties at $T\ll T_{c}$. The magnitude of this dip is strongly correlated with the RF surface resistance, impurity distribution near the surface, and $T_{\mathrm{c}}$. Additionally, we report measurements of the coherence peak in the AC conductivity of two Nb SRF cavities processed using distinct methods. By comparing recent theories developed to model this experimental data, we show that the frequency dip feature, larger coherence peak height, and reduction in the temperature-dependent surface resistance with RF current occur at minimal but finite levels of disorder.
@@ I Introduction
Superconducting radio frequency (SRF) niobium cavities are record-high quality factor $Q_{0}>10^{10}-10^{11}$ man-made resonators that serve as the primary accelerating structures in modern particle accelerators [1], longest coherence microwave quantum superconducting systems [2], and an ultrasensitive platform for dark sector and dark matter candidate searches [3, 4, 5, 6]. These ultra-high quality factors are enabled by the low microwave surface resistance of superconducting Nb, given by $R_{s}=G/Q_{0}$, where $G$ is a geometric factor. Typically, $R_{s}$ increases with the magnitude of the peak surface magnetic field $B_{p}$ due to increased screening currents. Extensive research over the past decades has allowed the discovery and mitigation of several such field-dependent physical mechanisms in various regimes [1]. This most recently includes proximity-coupled niobium nanohydrides [7] and their mitigation by oxygen inward diffusion [8] and the reduction of two-level system-driven dissipation by oxide removal [9, 2, 10].
@@
The AC complex conductivity enables further insight on quasiparticles and superconducting carriers. For frequencies far below the superconducting gap ($\hbar\omega\ll\Delta$), $\sigma_{1}$ exhibits the coherence peak, a nonmonotonic dependence of the quasiparticle conductivity with temperature that arises due to coherence factors in BCS theory [16] and is analogous to the Hebel-Slichter peak in nuclear spin relaxation [17]. This peak was originally demonstrated in niobium by Klein et al. [18] at 60 GHz, a frequency significantly higher than those typically used in modern RF accelerators. We show that the height of the coherence peak also correlates with the frequency dip in Fig.1.
In this article, we present a systematic study on a large number ($>$40) of bulk niobium SRF cavities revealing the anomalous decrease (dip) in cavity resonant frequency at temperatures just below $T_{\mathrm{c}}$ in nitrogen-doped cavities. In addition, we report that the other four (non-dip) characteristic temperature dependencies uniquely correspond to other explored surface treatments, and thus correlate strongly with cavity performance in $R_{T}$ with the stored field. The measured coherence peak in the complex conductivity of bulk niobium SRF cavities is fitted with a phenomenological model incorporating a pair-breaking term in the BCS [16] density of states. We comment on different models proposed by other groups which help provide insight on the origins of these various phenomena. Fits with theories suggest that the coherence peak, along with the dip phenomenon, high $Q_{0}$, and $R_{T}$ decrease with field, are signatures of improved superconducting properties and occur when extrinsic pair-breaking scattering and pair-conserving scattering from nonmagnetic impurities are minimal but finite.
@@ II Experimental Procedure
A large set of bulk niobium SRF cavities of TESLA [19] elliptical shape were employed for our studies. The resonant frequencies of these cavities varied from 650 MHz to 3.9 GHz, which allowed probing the low frequency limit of Nb ($\hbar\omega\ll\Delta$). All cavities went through a bulk electropolishing (EP) of 120 $\mu$m material removal and 800C hydrogen degassing step [20] followed by various surface treatments. These treatments included additional $\sim$40 $\mu$m EP, nitrogen infusion [21], 120C baking for 48 hours [22], 75/120C baking [23, 24] wherein cavities were first baked at 75C before the standard 120C bake, and nitrogen doping [11] with varying dopant concentrations. Dopant concentration is varied by altering the durations of baking in N atmosphere and in vacuum, which has the shorthand notation of minutes in N/minutes in vacuum (i.e., 2/6) and/or the amount of removal via EP.
All cavities were first cooled down to 2 K where standard measurements [25] of $Q_{0}$ as a function of the peak surface magnetic field $B_{\mathrm{peak}}$ have been performed; most cavities were tested again after cooling further down to $T<$ 1.5 K. To extract quasiparticle driven surface resistance $R_{T}$ from the measured $Q_{0}$, we used the methods laid out in [15]. Specifically, we subtracted off the residual resistance ($R_{\text{res}}=G/Q_{0}(<1.5\text{K})$) contribution from the total $R_{s}$ at 2 K, so that $R_{T}=R_{s}(2\text{K})-R_{\text{res}}$.
@@
To calculate the AC conductivity, we used the method utilized by Trunin et al. [29]. To extract the surface resistance, we measured $Q_{0}$ from 1.5 K up to 10 K using a combination of standard power balance measurements in continuous wave and decay modes along with VNA measurements and converted to surface resistance $R_{s}$. The surface reactance was measured utilizing the above discussed frequency vs temperature measurements from 6 K to 10 K and extrapolating to lower temperatures by means of fitting with the Halbritter routine [27] and using $X_{s}(T)=-2G\Delta f_{0}(T)/f_{0}+X_{n}$. We obtained the additive constant using the normal conducting surface resistance measured just above the transition temperature $R_{n}$(10 K). In the local limit, valid for dirty Nb, $R_{n}(\text{10 K})=X_{n}(\text{10 K})$. When the local limit is not applicable, as is true for high purity niobium [22], the correct relationship between $R_{s}$ and $X_{s}$ is obtained using the universal impedance curves plotted against the dimensionless parameter $\alpha$ calculated by Reuter and Sondheimer in the microwave region [30]. The parameter $\alpha$ is proportional to $\ell^{3}$/$\rho\ell$, where $\ell$ is the mean free path and $\rho\ell$ is a temperature independent material constant, which is 6 $\times$ 10-16 $\Omega$m2 for Nb [22]. Finally, the complex conductivity is calculated as
@@ III.3 Correlating $Q_{0}$ with $\Delta f_{dip}$
The evolution observed in Fig. 2 suggests correlation between the frequency dip and $R_{T}$; to investigate this, we plotted $R_{T}$ measured at 68 mT and 2 K against the dip magnitude for an ensemble of variously nitrogen doped cavities in Fig. 3. Indeed, we observe a linear trend, suggesting a potential relationship between the positive Q-slope and frequency dip phenomena. We note that the trend shown in Fig. 3 is relevant for the accelerator community as it facilitates a method for quickly estimating the temperature-dependent component of the surface resistance of N-doped cavities at high fields from low field frequency domain measurements, which may be practical when high power infrastructure is not readily available.
Figure 3: Temperature dependent surface resistance measured at 68 mT and a temperature of 2 K plotted against the dip magnitude for variously nitrogen doped cavities. Shown for reference are results obtained from a cavity post EP.
@@ III.5 Measurement of AC Complex Conductivity
To gain insight on the conditions under which different frequency features near $T_{c}$ occur, we measured the AC complex conductivity of two niobium cavities: i) the cavity presented in Fig. 2(a) with $\ell=550\pm 51$ nm which showed the standard feature near $T_{c}$ (which we call the EP cavity) and ii) a cavity post nitrogen doping which exhibited the prominent dip and $\ell=71\pm 3$ nm. We note that this latter mean free path is in agreement with the value obtained by extrapolating Fig. 2(c) down to $T_{c}=$ 9.04 K, yielding approximately 70 nm. Fig. 5(a) and its inset show the expected $Q_{0}(B_{p})$ and $R_{T}(B_{p})$ behaviors characteristic of cavities subjected to these surface treatment. Fig. 5(b) presents the real and the imaginary parts of the measure surface impedance of the two cavities.
@@
EP 3.70 $\times$ 10 -4 550 $\pm$ 51 36 $\pm$ 4 40 $\pm$ 0.1
N-Doped 5.88 $\times$ 10 -4 71 $\pm$ 3 57 $\pm$ 6 48 $\pm$ 0.4
@@
Figure 5: (a) Quality factor vs peak magnetic field of two 1.3 GHz cavities subjected to EP or N-doping. Inset shows the temperature dependent surface resistance at 2 K as a function of peak magnetic field. (b) Surface reactance of the EP and N-doped cavities as a function of temperature; inset depicts the surface resistance as a function of temperature. (c) Real ($\sigma_{1}$) and imaginary ($\sigma_{2}$) components of the AC conductivity normalized to the normal conducting value $\sigma_{n}$. Error bars show the uncertainty in the calculated values (see text for details). Dashed and solid lines show fits.
The calculated AC complex conductivities of the N-doped and EP cavities are presented in Fig. 5(c). The error bars indicate measurement uncertainty. For the N-doped cavity, this uncertainty is primarily due to the measurement process itself, whereas for the EP cavity, the dominant uncertainty comes from the fitted value of $\ell$. We note that as the technique used to measure the surface impedance sums over the inner RF surface, these curves represent the average cavity response.
The non-monotonic dependence in the real part of the complex conductivity in BCS superconductors, called the coherence peak, arises due to the singularity in the superconducting density of states at the gap edge [16, 34]. According to Dynes et al. [35], inelastic scattering smears the singularity and introduces subgap quasiparticle states. By incorporating the Dynes smearing parameter $\Gamma$ into a modified version of the Mattis-Bardeen conductivity, it is possible to obtain a phenolomenological measure of the pair-breaking processes present in a system by fitting the coherence peak using
@@
From Fig. 5(c), we observe that both the EP and N-doped cavities exhibit the so-called coherence peak, with the maxima at $\sim$0.9 $T/T_{c}$ differing in amplitude. This suggests a variation in the quasiparticle behavior within the RF layer due to distinct impurity distributions. The stark difference observed in $\sigma_{2}/\sigma_{n}$ between the two cavities highlights a variance in the superconducting condensate behavior.
The dashed and solid curves in Fig. 5(c) are calculated using Eq. 2-4. We find that both $\sigma_{1}/\sigma_{n}$ and $\sigma_{2}/\sigma_{n}$ of the EP cavity are best modelled with an average superconducting gap $\Delta_{0}=1.5$ meV and an inelastic scattering parameter $\Gamma/\Delta_{0}$ = 0.025. The N-doped cavity is instead better fitted with a larger average superconducting gap $\Delta_{0}=1.6$ meV and a lower level of pair-breaking ($\Gamma/\Delta_{0}$ = 0.0066). Note that a larger coherence peak occurs for the cavity with greater elastic scattering (shorter $\ell$) due to the presence of N interstitial. According to Anderson’s theorem [36], elastic scattering off nonmagnetic impurities has no effect on superconductivity. As a result, while N impurities increase scattering, they likely mitigate depairing processes which results in a larger coherence peak. Notably, when the fit parameters from the N-doped data are applied to Eq. 2-4 at 60 GHz, the calculated $\sigma_{1}/\sigma_{n}$ results roughly match Klein’s local-limit measurements [18].
@@ IV Discussion
The frequency phenomena and trends observed in this work have stimulated the development of new theoretical approaches which seek to provide insight into their microscopic origins. Ueki et al. provided the first explanation of the anomalous near $T_{c}$ frequency phenomena [37] by calculating the complex surface impedance via Slater’s approach to solving Maxwell’s equations for the case of a cavity while considering anisotropy in the superconducting gap [32] and inhomogeneous disorder within the penetration depth. By varying the level of disorder within the cavity screening region, Ueki calculated the foot, dip+bump, and dip features in the frequency response just before $T_{c}$. Moreover, the authors obtained the predicted cavity $Q_{0}$ as a function of disorder. We note that the frequency dip feature and a maximum in $Q_{0}$ were obtained for intermediate levels of disorder. Ueki also correctly reproduced the frequency dependence of the dip feature reported in Fig. 4. Zarea et al, instead, derived the eigenvalue equation for the fundamental TM mode of a cylindrical RF cavity which considers the current response obtained with the Keldysh formulation of the quasiclassical theory of superconductivity [38]. The result considers the confinement of the normal conducting and superconducting currents within the RF layer and shows that the anomalous frequency dip is a product of the competition between the normal metal skin depth and the London penetration depth. An increase in impurity concentration not only affects the penetration depth but also the classification of superconducting niobium from a weak type-I to a weak type-II superconductor [39]. As a result, the presence of interstitial nitrogen can influence this competition between length scales. We note that Zarea also found that the frequency dip and a maximum in cavity $Q_{0}$ occur at intermediate levels of disorder.
@@
The results presented in Fig. 5 have also stimulated theoretical work by other groups which further improves our understanding of the mechanisms behind the improved performance of niobium SRF cavities in the presence of various near-surface impurity profiles. Herman and Hlubina [40] used the Nam approach [41] to calculate the optical response of a superconductor in the limit where energy tends to zero and with pair-conserving and pair-breaking elastic-scattering processes built into the energy and gap functions. The result describes the effect of subgap quasiparticle states on the optical conductivity and utilizes the concept of “Dynes superconductors.” The authors present excellent fits to the real and imaginary parts of the AC complex conductivity of both cavities presented in Fig. 5 and find that pair-breaking processes dominate the height of the coherence peak in $\sigma_{1}/\sigma_{n}$. Herman and Hlubina also highlight that the ratio between pair-breaking and pair-conserving scattering rates is key in determining the coherence peak extent and show that cleaner superconductors counterintuitively yield lower peak heights. Kubo, instead, used the Dynes $\Gamma$ in the Eilenberger formalism of the BCS theory to calculate several superconductor quantities in the presence of varying levels of nonmagnetic impurity scattering rates as a function of temperature and energy [42]. These calculations also well model the results presented in Fig. 5 and, similar to Herman and Hlubina, show that the ratio between pair-breaking and nonmagnetic impurity scattering rates is key in determining the height of the coherence peak. Moreover, Kubo finds that there exists a minimum in the surface resistance at finite levels of pair-breaking and pair-conserving scattering. We note that the phenomenologically derived Eq. 2-4 in the present manuscript reproduce the salient feature of the above theories that greater smearing of the density of states yields a reduction of the coherence peak height and the zero-temperature value of $\sigma_{2}/\sigma_{n}$.